OpenStudy (anonymous):

How many pounds of nuts worth $1.10 per pound must be mixed with 36 pounds of candy worth $1.60 per pound to produce a party mix worth $1.50 per pound?

OpenStudy (anonymous):

lets put the number of pounds to be added as \(x\) at the moment you have 36 pounds, so if you add \(x\) pounds you will have \(36+x\) pounds the 36 pounds are worth \(36\times 1.6=57.6\) and adding \(x\) pounds will give you a total price of \(57.6+1.1x\) which you want to be \(1.5\times (36+x)\) so set \[57.6+1.1x=1.5(36+x)\] and solve for \(x\)

OpenStudy (anonymous):

probably easier to get rid of the annoying decimal by multiplying both sides by 10 and get \[576+11x=15(36+x)\]

OpenStudy (anonymous):

hope it is routine from there, if not let me know